suppose a factory contains two machines. there is a 19 % probability that the first machine will break down and a 22 % chance the second machine breaks down. the two machines are independent.
A. what is probability both machines break down
B. what is the probability AT LEAST one machine breaks down
C. What is probability that only one breaks down.
A. Pr (A and B) = .0418
B. ?
C. ?

Answer by stanbon(75887)

suppose a factory contains two machines. there is a 19 % probability that the first machine will break down and a 22 % chance the second machine breaks down. the two machines are independent.
A. probability both machines break down = (0.19)(0.22)
------------------------------------
B. probability AT LEAST one machine breaks down = 1-(p(neither breaks down)
= 1 - [0.81*0.78]
---------------------------------
C. probability that only one breaks down = 1 - P[both break down]-P[neither
breaks down] = 1 - (ans. to A)- 0.81*0.78
================================================
Cheers,
Stan H.
================================================
A. Pr (A and B) = .0418
B. ?
C. ?